# Backtracking New Q-Newton’s method, Newton’s flow, Voronoi’s diagram and Stochastic root finding

### 02 Jan 2024 v1

@@ 1. Introduction
After presentations of BNQN in ICIAM 2023 (Waseda University) and a satellite conference, a couple of candidates, which have a lot of connections in mathematics and physics, for explaining the pictures obtained in [19] have emerged:
@@
In this paper, we will explore in depth the possible links mentioned above between BNQN, Newton’s flow and Voronoi’s diagrams, when finding roots of meromorphic functions. We verify with many different geometric configurations of the roots, that indeed the link is very apparent and well above a random coincidence. We work with three different versions of Newton’s flow, with two versions of Newton’s method, as well as Random Relaxed Newton’s method. For the readers’s convenience, we will recall in enough detail on Newton’s method and Random Relaxed Newton’s method in the next Section. It seems that the picture one obtains from BNQN is a better version of what one obtains with Newton’s method v2, and while BNQN is not a flow method it somehow attains many good features of flow methods. From these experiments, another interesting new phenomenon seems to occur when one applies Newton’s flow to $f/f^{\prime}$, even if one only wants to find roots of $f$ and even if $f$ has only roots of multiplicity $1$.
@@ 2.3.3. A main theoretical result for finding roots of meromorphic functions by BNQN
2) If $z^{*}$ is a cluster point of $\{z_{n}\}$ (that is, if it is the limit of a subsequence of $\{z_{n}\}$), then $z^{*}$ is a critical point of $F$.
@@ 2.5. Voronoi’s diagrams
There are many versions of Voronoi’s diagrams. In this paper, we consider the most basis version, that of a finite set of points $z_{1},\ldots,z_{m}$ in $\mathbf{R}^{2}$, with the Euclidean distance $d_{E}(.,.)$. If $z\in\mathbf{R}^{2}$ is such that there is a unique $i\in\{1,\ldots,m\}$ for which $d_{E}(z,z_{i})>\max_{j\not=i}d_{E}(z,z_{j})$, then we say that $z$ belongs to the Voronoi’s cell of the point $z_{i}$. The boundary consists of lines where the maximum of $\max_{j}d_{E}(z,z_{j})$ is attained at two or more points $z_{i}$’s.
@@ 3.1. Settings
3) Colours and interpretations of the results: If the factorisation of a function $f$ is written as (note that the order $z_{1}$, $z_{2}$, $\ldots$ is important) $f(z)=(z-z_{1})^{n_{1}}(z-z_{2})^{n_{2}}\ldots$, then we will colour the basins of attraction as follows: that for the first root has the green colour, that for the second root has the yellow colour, that for the third root has the blue colour, that for the fourth root has the red colour, that for the fifth root has the pink colour, that for the sixth root has the cyan colour, that for the seventh root has the orange colour, that for the eighth root has the purple colour. Other points (e.g. points whose corresponding sequence goes to infinity or does not converge) will have the black colour.
@@ 3.2.2. Transcendental functions
Figure 26. Basins of attraction for finding roots of the function $f_{7}e^{z}$ by different methods. Pictures are referenced to from top to bottom, from left to right. Row 1: left picture is for Newton’s method, right picture is for Random Relaxed Newton’s method. Row 2: left picture is for BNQN, right picture is for BNQN v2. Row 3: left picture is for Newton’s flow, right picture is for BNQN v2. Row 4: Voronoi’s diagram. The method Newton’s method vOptimization encounters errors, while Newton’s flow vOptimization takes too long time to finish.
Figure 27. Basins of attraction for finding roots of the function $f_{17}e^{z}$ by different methods. Pictures are referenced to from top to bottom, from left to right. Row 1: left picture is for Newton’s method, right picture is for Random Relaxed Newton’s method. Row 2: left picture is for BNQN, right picture is for BNQN v2. Row 3: left picture is for Newton’s flow, right picture is for BNQN v2. The method Newton’s method vOptimization encounters errors, while Newton’s flow vOptimization takes too long time to finish.
@@ 4. Conclusions and further directions
$(f/f^{\prime})^{\prime}=\frac{ff^{\prime}}{(f^{\prime})^{2}-ff"},$

### 08 Jan 2024 v2

@@ 1. Introduction
After presentations of BNQN at ICIAM 2023 (Waseda University) and a satellite conference, a couple of candidates, which have a lot of connections in mathematics and physics, for explaining the pictures obtained in [19] have emerged:
@@
In this paper, we will explore in depth the possible links mentioned above between BNQN, Newton’s flow and Voronoi’s diagrams, when finding roots of meromorphic functions. We verify with many different geometric configurations of the roots, that indeed the link is very apparent and well above a random coincidence. We work with three different versions of Newton’s flow, with two versions of Newton’s method, as well as Random Relaxed Newton’s method. For the readers’s convenience, we will recall in enough detail on Newton’s method and Random Relaxed Newton’s method in the next Section. It seems that the picture one obtains from BNQN is a better version of what one obtains with Newton’s method for Optimization, and while BNQN is not a flow method it somehow attains many good features of flow methods. From these experiments, another interesting new phenomenon seems to occur when one applies Newton’s flow to $f/f^{\prime}$, even if one only wants to find roots of $f$ and even if $f$ has only roots of multiplicity $1$.
@@ 2.3.3. A main theoretical result for finding roots of meromorphic functions by BNQN
2) If $z^{*}$ is a cluster point of $\{z_{n}\}$ (that is, if it is the limit of a subsequence of $\{z_{n}\}$), then $z^{*}$ is a critical point of $F$. Moreover, in this case the whole sequence $\{z_{n}\}$ converges to $z^{*}$.
@@ 2.5. Voronoi’s diagrams
There are many versions of Voronoi’s diagrams. In this paper, we consider the most basis version, that of a finite set of points $z_{1},\ldots,z_{m}$ in $\mathbf{R}^{2}$, with the Euclidean distance $d_{E}(.,.)$. If $z\in\mathbf{R}^{2}$ is such that there is a unique $i\in\{1,\ldots,m\}$ for which $d_{E}(z,z_{i})<\min_{j\not=i}d_{E}(z,z_{j})$, then we say that $z$ belongs to the Voronoi’s cell of the point $z_{i}$. The boundary consists of line intervals where the maximum of $\min_{j}d_{E}(z,z_{j})$ is attained at two or more points $z_{i}$’s.
@@ 3.1. Settings
3) Colours and interpretations of the results: If the factorisation of a function $f$ is written as (note that the order $z_{1}$, $z_{2}$, $\ldots$ is important) $f(z)=(z-z_{1})^{n_{1}}(z-z_{2})^{n_{2}}\ldots$, then we will colour the basins of attraction as follows: that for the first root has the green colour, that for the second root has the yellow colour, that for the third root has the blue colour, that for the fourth root has the red colour, that for the fifth root has the pink colour, that for the sixth root has the cyan colour, that for the seventh root has the orange colour, that for the eighth root has the purple colour. Other points (e.g. points whose corresponding sequence goes to infinity or does not converge to one of the roots) will have the black colour.
@@ 3.2.2. Transcendental functions
Figure 26. Basins of attraction for finding roots of the function $f_{7}e^{z}$ by different methods. Pictures are referenced to from top to bottom, from left to right. Row 1: left picture is for Newton’s method, right picture is for Random Relaxed Newton’s method. Row 2: left picture is for BNQN, right picture is for BNQN v2. Row 3: left picture is for Newton’s flow, right picture is for Newton’s flow vFraction. Row 4: Voronoi’s diagram. The method Newton’s method vOptimization encounters errors, while Newton’s flow vOptimization takes too long time to finish.
Figure 27. Basins of attraction for finding roots of the function $f_{17}e^{z}$ by different methods. Pictures are referenced to from top to bottom, from left to right. Row 1: left picture is for Newton’s method, right picture is for Random Relaxed Newton’s method. Row 2: left picture is for BNQN, right picture is for BNQN v2. Row 3: left picture is for Newton’s flow, right picture is for Newton’s flow vFraction. The method Newton’s method vOptimization encounters errors, while Newton’s flow vOptimization takes too long time to finish.
@@ 4. Conclusions and further directions
$(f/f^{\prime})/(f/f^{\prime})^{\prime}=\frac{ff^{\prime}}{(f^{\prime})^{2}-ff"},$